OpenStudy (anonymous):

what is the log of √24.3 ?

OpenStudy (lgbasallote):

are you asked to expand it? or the exact value?

OpenStudy (anonymous):

exact value

OpenStudy (lgbasallote):

okay..here's a hint: \[\large \log \sqrt{24.3} \implies \log_{10} (\sqrt{24.3}) \implies \log \sqrt{24.3} \div \log 10\] does that help?

OpenStudy (anonymous):

it is 0.692803

OpenStudy (lgbasallote):

or actually...you can just input in a calculator \[\large \log \sqrt{24.3}\]it will automatically give you the answer

OpenStudy (anonymous):

Thanks a lot both :)

OpenStudy (anonymous):

i hv done the same way .....@igbasallote...lolz

OpenStudy (rishabh.mission):

0.692803